OpenStudy (anonymous):

5x - 4y = 3 2x+ 7y = -10, use the elimination method help solve

OpenStudy (anonymous):

5x - 4y = 3 2x+ 7y = -10, use the elimination method help solve

OpenStudy (anonymous):

2(5x-4y)=2(3) 5(2x+7y)=5(-10) 2(5x-4y)=2(3) - 5(2x+7y)=5(-10) ----------------- 0x -43y = 56 y= -56/43 5x-4(-56/43)=3 5x+224/43=2 5x=129/43-224/43 5x=-95/43 x=-19/43